2 particles p and q strts moving from A  with velocity 15m/s and 20m/s respectively .2 particles moves with the acelaration equal in magnitude but opposite in direction P overtakes Q at the point B then its velocity 30 m/s .velocity of Q at point B is 
 

Grade:12

1 Answers

Susmita
5 years ago
P overtakes Q means P has positive acceleration and Q has negative acceleration.
Let P take a time t to overtake Q.Means P and Q both run for time t before they meet at some point.
P starts with initial speed u=15m/s.At the time of crossing P's speed is v=30m/s.
v=u+ft
Or,30=15+ft
Or,ft=15
Q starts with a speed u=20 m/s and acceleration -f.So its velocity at the time of crossing is
v=u-ft=20-ft=20-15=5 m/s
If it helps please approve it.